### v3

# UCFE: A User-Centric Financial Expertise Benchmark for Large Language Models

Abstract
This paper introduces the UCFE: User-Centric Financial Expertise benchmark, an innovative framework designed to evaluate the ability of large language models (LLMs) to handle complex real-world financial tasks. UCFE benchmark adopts a hybrid approach that combines human expert evaluations with dynamic, task-specific interactions to simulate the complexities of evolving financial scenarios. Firstly, we conducted a user study involving 804 participants, collecting their feedback on financial tasks. Secondly, based on this feedback, we created our dataset that encompasses a wide range of user intents and interactions. This dataset serves as the foundation for benchmarking 11 LLMs services using the LLM-as-Judge methodology. Our results show a significant alignment between benchmark scores and human preferences, with a Pearson correlation coefficient of 0.78, confirming the effectiveness of the UCFE dataset and our evaluation approach. UCFE benchmark not only reveals the potential of LLMs in the financial domain but also provides a robust framework for assessing their performance and user satisfaction.
UCFE: A User-Centric Financial Expertise Benchmark for Large Language Models
Yuzhe Yang1††thanks: Equal contribution: yuzheyang@link.cuhk.edu.cn, yf_zhang@smail.nju.edu.cn, huyan@cuhk.edu.cn, Yifei Zhang211footnotemark: 1, Yan Hu111footnotemark: 1, Yilin Guo1, Ruoli Gan1, Yueru He3, Mingcong Lei1, Xiao Zhang3, Haining Wang2, Qianqian Xie322footnotemark: 2, Jimin Huang3, Honghai Yu222footnotemark: 2, Benyou Wang1††thanks: Corresponding authors: qianqian.xie@thefin.ai, hhyu@nju.edu.cn, benyouwang@cuhk.edu.cn 1The Chinese University of Hong Kong, Shenzhen, 2Nanjing University, 3The Fin AI https://github.com/TobyYang7/UCFE-Benchmark
1 Introduction
Recent advances in large language models (LLMs) have expanded their potential applications in finance Wu et al. (2023); Huang et al. (2023); Kim et al. (2024). Finance professionals are increasingly using LLMs to solve specialized financial tasks Li et al. (2023a); Xie et al. (2024); Yang et al. (2023); Zhang et al. (2023b), including explorations into LLM-powered financial agents Li et al. (2024); Yang et al. (2025). Financial tasks are inherently complex, involving specialized context, financial terminologies, legal intricacies, and dynamic markets that involve information with high noise-to-signal ratio Pagano (1993); Mullainathan and Spiess (2017); Li et al. (2018), which adds significant challenges for LLMs to address. Accurate analysis of financial information is crucial, as even minor ignorance in a signal or market information can lead to substantial financial losses Tversky and Kahneman (1981); Thaler (2008); Mohamed et al. (2024).
To be effective, LLMs need to swiftly adapt to fiscal policy changes, market fluctuation, extreme events, and global factors, identifying key signals within real-time data to manage volatility and mitigate risks Gueta et al. (2024); Yadav et al. (2024). Financial markets can react instantly to news, making it crucial for LLMs to process information in near real-time by rapidly consolidating unstructured, real-time data from multiple sources Nguyen and Tulabandhula (2023); Tong et al. (2024). Despite LLMs’ improving accuracy on tasks like sentiment analysis, market prediction, and risk assessment Wimmer and Rekabsaz (2023); Lopez-Lira and Tang (2023); Rizinski et al. (2024), these models still face significant limitations, such as their reliance on static datasets and challenges in handling real-time data, which hinders their real-world applicability in dynamic financial contexts. Moreover, the evolving nature of financial regulation adds another layer of complexity, requiring LLMs to continuously update their knowledge to remain compliant and useful Yao et al. (2024); He et al. (2024). These limitations highlight the need for a more dynamic evaluation framework that assesses LLMs’ performance under real-time, evolving financial conditions, ensuring they can handle not only static tasks but also the unpredictable nature of real-world financial environments.
To address these challenges in financial domain, we propose a novel framework, the User-Centric Financial Expertise Benchmark, designed to evaluate the ability of LLMs to handle financial tasks in real-world scenarios. Figure 1 provides an overview of the framework. The UCFE benchmark has the following key features:
User-Centric Design: Based on preliminary surveys and research, we categorized the target user group into four distinct types: analysts, financial professionals, regulatory professionals, and the general public. Using questionnaires in Appendix A, we gathered insights into the primary needs and practical applications of each group. This allowed us to refine the user categories for more targeted evaluations, in which LLMs are prompted to simulate roles representative of each group. We developed 17 task types tailored to these user profiles, encompassing 330 data points that include multi-round dialogues in both zero-shot and few-shot settings. More details of the dataset will be explained in Section 4.
Dynamic Interactions: In the few-shot tasks, each user group follows a task-oriented approach, where an initial action goal is defined. Users articulate their professional needs and specific task requirements through successive interactions. To simulate real-world user scenarios, we employ authentic datasets that closely mirror actual financial scenarios. This dynamic interaction setup ensures that LLMs are not only responding to isolated queries but are also engaging in an evolving dialogue, adjusting their responses based on the ongoing professional needs expressed by the user. This method provides a more accurate reflection of how LLMs would perform in practical, task-specific financial contexts.
In summary, this work makes the following key contributions: (1) We propose a new framework that combines human expert judgments with LLMs to assess their ability to handle increasingly complex financial tasks. (2) By leveraging dynamic, user-centric interactions, this work probes the boundaries of LLM capabilities by examining how well LLMs adapt to evolving professional needs and increasingly complex task requirements, which provides a deeper understanding of their potential and limitations in addressing real-world financial scenarios.
2 Related Work
2.1 Financial Benchmark
FLARE Xie et al. (2023) evaluates models on five financial tasks 111sentiment analysis, news headline classification, named entity recognition, question answering, and stock movement prediction. Existing benchmarks Zhang et al. (2023a); Li et al. (2023b); Yuan et al. (2024) primarily use multiple-choice questions to assess models’ domain knowledge, with questions sourced from real-world financial documents and publicly available financial reports or websites, covering a wide range of topics such as finance, economy, accounting, and certification. MMMU and MMMU-PRO Yue et al. (2024a, b) extend beyond traditional financial NLP tasks by incorporating multimodal inputs to better evaluate models in more complex financial tasks. Although these benchmarks have advanced the evaluation of financial language models, they predominantly consider structured NLP tasks with deterministic answers and rely heavily on multiple-choice questions or tasks with specific answers, such as sentiment analysis and named entity recognition. This limits their ability to assess generative capabilities, which are essential for simulating real-world financial applications Krause (2023); Koa et al. (2024).
2.2 User-Centric Framework
The implementation of user-centric models involves integrating users into core business processes to harness their creative potential, which has been successfully demonstrated by companies like LEGO, IBM, and Coloplast Hienerth et al. (2011); Kwon et al. (2021). EUCA framework, a practical prototyping toolkit designed to make AI systems explainable to non-technical end-users, provides twelve end-user-friendly explanatory forms that do not require technical knowledge to bridge the gap between technical creators and non-technical users Jin et al. (2021). In the financial domain, user-centric explainability is also crucial in algorithmic decision-making systems like robo-advisors Naveed et al. (2022); Roveda et al. (2023); Pisoni and Díaz-Rodríguez (2023). Research has highlighted the importance of providing transparent and comprehensible explanations to users, which indicated that user trust and confidence in financial applications are positively correlated with the transparency and comprehensibility of the explanations provided How et al. (2020); Deo and Sontakke (2021); Xu et al. (2024).
3 Background
Recent advancements in LLMs have demonstrated significant potential in addressing complex financial tasks. Numerous organizations are now actively training their own LLMs, aiming to enhance their performance by incorporating extensive domain-specific knowledge. For instance, FinGPT Yang et al. (2023), which applies supervised fine-tuning to the LLaMA model, has shown notable improvements in financial tasks. Through continued pretraining, models like FinLLaMA Xie et al. (2024) have further advanced LLM performance across various metrics. These developments highlight the growing demand and potential for LLMs in the financial domain, both in academia and industry.
From a technical standpoint, LLMs have undoubtedly reduced costs and improved efficiency by quickly processing vast amounts of financial text data with commendable performance. However, the challenges of developing real-world financial applications extend beyond technical issues. These challenges include business requirements, industry-specific barriers, data privacy concerns, accountability, and ethical considerations Nie et al. (2024); Yao et al. (2024), along with a gap in understanding between LLMs, functioning as AI assistants, and the specific needs of financial experts.
As discussed in Section 2.1, existing benchmarks largely focus on technical metrics such as accuracy and efficiency, often ignoring these broader challenges. By emphasizing only technical aspects, such benchmarks fail to address the real-world complexities of financial applications, where business rules, legal frameworks, and human judgment play crucial roles. This makes non-technical aspects, particularly human-AI interaction in finance, comparatively under-explored. Human-AI interaction is critical in financial settings, as it affects decision-making, user trust, and the effective integration of AI systems into the financial workflow. Without considering these factors, current benchmarks offer an incomplete evaluation, limiting the practical relevance of LLMs for real-world financial applications. There is a pressing need for evaluation frameworks that not only assess technical performance but also account for the nuanced interplay between AI systems and financial professionals in complex environments.
In addition to these challenges, the rise of FinTech companies such as Robinhood 222https://robinhood.com/us/en/ has spurred increasing public interest in finance and trading. A growing number of individuals, many without formal financial education, are seeking accessible ways to manage their finances and participate in the market. For these users, LLMs have become a significant source of financial learning and advice. However, the accuracy, flexibility, and contextual understanding of LLMs are crucial to meet this emerging demand. Models must not only provide correct and relevant information but also adapt to the diverse financial needs of the general public, from basic education to advanced financial management strategies.
In this context, the UCFE benchmark will mainly focus on evaluating the interaction between LLMs and humans, as the improvements in user experience often have greater practical significance than gains in task-specific metrics. By introducing this new framework, we aim to offer deeper insights into the future development of financial LLMs, aligning model performance more closely with human preferences across multiple dimensions. This approach is intended to provide a more holistic understanding of how LLMs can better serve real-world financial applications, ultimately leading to more user-centric AI solutions.
4 User-Centric Financial Expertise Dataset
4.1 User Preference Alignment
To align our dataset more closely with real-world financial tasks and user needs, we conducted a survey to gather insights into how users engage with financial scenarios. Participants completed a questionnaire designed to capture key aspects of their interactions with financial tasks, focusing on their roles, levels of experience, and the types of tasks they typically perform. The survey included questions about participants’ familiarity with financial analysis, preferred sources of information, and their engagement preferences regarding financial tasks.
Feedback was solicited from participants across three main areas:
Participant Demographics: Information on the participants’ backgrounds and expertise.
Detailed Interaction with Financial Tasks: Assessment of participants’ experiences and interactions with specific financial tasks.
Financial Scenario Coverage: Evaluation of how well the tasks reflected real-world financial scenarios.
The full questionnaire can be found in Appendix A.
4.2 Dataset Construction
| User | Familiarity | Importance | |
| Total | 804 | 458 | 660 |
| Student (Finance-related) | 167 | 148 | 155 |
| Financial Professional | 83 | 83 | 83 |
| Regulatory Professional | 51 | 47 | 50 |
| General Public | 136 | 49 | 82 |
| Non-Finance Professional | 87 | 37 | 70 |
| Student (Non-finance) | 208 | 79 | 163 |
| Other | 72 | 15 | 57 |
Based on the results of the survey shown in Table 1, we recognized the necessity of constructing a multi-round finance dialogue benchmark that serves both the finance-related and non-finance groups. The survey revealed diverse user intentions and varying levels of financial expertise, underscoring the need for a benchmark that can accommodate a broad spectrum of scenarios. By catering to these different groups, we aimed to capture a comprehensive range of dialogue interactions, from complex financial analysis to simpler, more general financial inquiries, ensuring the dataset reflects real-world variations in user needs and knowledge levels.
To establish the multi-round dialogue benchmark, we meticulously selected sources that encompass authoritative financial reports, regulatory documents, and accessible online resources based on the survey results shown in Appendix B.2. This selection process was designed to ensure that the dataset meets both the technical demands of financial professionals and the practical needs of general users. By synthesizing insights from diverse user experiences and expert evaluations, we aimed at creating a dataset that facilitates effective multi-round interactions, ultimately enhancing the user experience in financial analysis.
4.3 Tasks
Table 2 and 3 provide the statistical breakdown of the UCFE benchmark, with all data sourced from the previous user survey targeting various user groups. The UCFE benchmark encompasses both few-shot and zero-shot tasks, with a total of 17 distinct tasks covering a broad range of financial scenarios. These tasks are specifically designed to reflect practical financial needs, including but not limited to market information summarization, asset valuation, and regulatory compliance assessments. The multi-turn nature of these tasks emphasizes dynamic user interaction and adaptive decision-making.
Figure 2 presents the visualization of the 25 most common root verbs (inner circle) and their top 4 associated direct noun objects (outer circle), providing insights into the types of financial interactions covered by the dataset. The diversity of verb-noun pairs highlights the wide range of financial operations and decision-making processes represented, ensuring the benchmark tasks reflect the complex and varied language used in financial contexts.
In addition, Figure 3 shows the distribution of input lengths for both test and evaluation queries, revealing significant variance in task complexity. Shorter queries require concise outputs, while longer inputs demand deeper comprehension and detailed responses. This variance challenges models not only to generalize across different task types but also to adapt their performance based on the complexity of the input, making it essential for evaluating LLMs’ scalability and versatility in real-world financial tasks.
| Category | Task | Source | Target User Group |
| Few-shot | Analyst Simulation | TCL Annual Report & Analyst Report | Senior Analyst |
| Asset Valuation Reporting | EastMoney | Analyst | |
| Company Evaluation Reporting | Analyst Report | Analyst | |
| Corporate Operation Analysis | Analyst Report | Analyst | |
| Credit Risk Evaluation | GPT-4 Generated | Analyst | |
| Financial Knowledge Consulting | Investopedia1 | General Public & Financial Professional | |
| Financial Regulation Consulting | Securities Law2 | General Public & Financial Professional & Regulatory Professional | |
| Industry Report Summarization | EastMoney | General Public & Financial Professional | |
| Insider Trading Detection | Securities Regulatory Commission3 | Regulatory Professional | |
| Investment Strategy Evaluation | Seeking Alpha4 | Analyst | |
| Investment Strategy Optimization | Financestrategists5 | Analyst | |
| Newshare Evaluation Reporting | Stock.us6 | Analyst | |
| Prospectus Risk Summarization | Prospectus & Inquiry Letter7 | General Public & Financial Professional | |
| Zero-shot | Stock Price Prediction | A-stock Statistics | General Public & Financial Professional |
| Negative Information Detection | EastMoney | General Public & Financial Professional | |
| Financial Indicator Calculation | CPA & CFA | General Public & Financial Professional | |
| Financial Text Summarization | News Headlines | General Public & Financial Professional |
- 1
- 2
- 3
- 4
- 5
- 6
- 7
| Task Type | Number of Tasks | Number of Questions |
| Zero-shot Tasks | 4 | 80 |
| Few-shot Tasks | 13 | 250 |
| Total | 17 | 330 |
5 UCFE Benchmark
In this section, we provide an overview of the technical details and evaluation pipeline of the UCFE benchmark. As shown in Figure 4, the evaluation starts by selecting finance-specific tasks (introduced in Section 4.3), where the model acts as an AI assistant. GPT-4o333https://openai.com/index/gpt-4o-system-card/ simulates user interactions, generating dialogue data based on realistic behavior. Using LLMs to simulate user roles is common in recent research Inaba et al. (2024). To minimize model bias, we established evaluation criteria (detailed in Section 5.2). Model outputs are then compared in pairs, with Claude-3.5-Sonnet444https://www.anthropic.com/news/claude-3-5-sonnet as the evaluator, following the common practice LLm-as-judge framework for evaluation Liu et al. (2023). Each model’s performance is measured using Elo scores, which offer a relative comparison of capabilities. Finally, results are compared against human expert preferences to ensure evaluation robustness.
5.1 Evaluation Method
We use the Elo rating system for model evaluation, which is well-suited for comparing multiple models. This system is widely applied in competitive environments, such as match result prediction in sports like association football Hvattum and Arntzen (2010); Chiang et al. (2024). Its key advantages are:
Dynamic Adjustments: Elo ratings are continuously updated based on relative model performance, making it ideal for frequent comparisons.
Scalability & Efficiency: New models can be added without retesting all previous ones, saving time and API costs.
Each model starts with an Elo rating of , which is updated after every comparison task. For each task, dialogues generated by the target model and the base model are compared using specific prompts. A Claude-based model evaluates the comparison as a win, loss, or tie, and the Elo ratings are updated using the formula:
where is the updated rating, is the current rating, is the result (1 for a win, 0.5 for a tie, and 0 for a loss), and is the expected result, computed as:
Here, is the opponent’s rating, is set to 400, and is 4. These parameters control the magnitude of rating updates. This process repeats for each task, and the final Elo ratings reflect the models’ comparative performance across all tasks.
5.2 Experimental Settings
In the experiments, GPT-4o is used as the user simulator to generate queries and simulate real-world conversations. Claude-3.5-Sonnet serves as the evaluator to compare model responses, ensuring a clear separation between testing and evaluation to minimize bias.
For dialogue simulations, the temperature is set to 0.5 with no token limit. We tested financial-specific LLMs (7B to 70B parameters) along with their backbone models and included general-purpose models like GPT-4o and GPT-4o-mini, accessed via APIs. Table 4 lists all models used.
To mitigate positional bias in LLM evaluations Li et al. (2023c), we shuffled the input order during dialogue comparisons. To further minimize evaluator bias, such as misinformation or cognitive bias Talboy and Fuller (2023), we designed the evaluation prompts based on two key criteria:
Source Information Content: Categorized into Answer, Must Contain, At Least Contain, Should Contain, Encourage Contain, and Forbid Contain, guiding LLMs to make accurate content-based choices.
Evaluation Standards: Focused on User Needs Fulfillment, Fact Accuracy, Responsibility & Fairness, Response Depth & Richness, Hallucination, and Conciseness & Clarity, ensuring a thorough assessment.
The full evaluation prompt is available in Appendix E.
| Model | Type |
| CFGPT2-7B1 Li et al. (2023a) | Financial |
| GPT-4o | General |
| GPT-4o-mini | General |
| InternLM2.5-7B-Chat Cai et al. (2024) | General |
| Llama-3.1-70B-Instruct AI@Meta (2024) | General |
| Llama-3.1-8B-Instruct | General |
| Llama3-XuanYuan3-70B-Chat Zhang et al. (2023b) | Financial |
| Palmyra-Fin-70B-32k team (2024) | Financial |
| Qwen2.5-14B-Instruct Team (2024) | General |
| Qwen2.5-7B-Instruct | General |
| Tongyi-Finance-14B-Chat2 | Financial |
1
The backbone model of CFGPT2-7B is InternLM2-7B.
2
The backbone model of Tongyi-Finance-14B-Chat is Qwen-14B.
5.3 Overall Results
Table 5 presents a comprehensive overview of model performance across the 17 distinct financial tasks within the UCFE benchmark. A key finding is the consistent outperformance of financially-specialized LLMs (Tongyi-Finance-14B-Chat, CFGPT2-7B, and Palmyra-Fin-70B-32k) compared to original backbone models ( Qwen series, InternLM series and Llama series). This performance gap is not uniform across all tasks, highlighting the varied strengths and weaknesses of different model architectures and training strategies when applied to the complexities of the financial domain. While general-purpose models demonstrate a baseline level of competence, their performance often lags significantly on tasks requiring in-depth financial knowledge, precise terminology, and adherence to regulatory constraints.
To ensure the robustness of our evaluation and mitigate potential biases inherent in using a single LLM as a judge (Claude-3.5-Sonnet), we conducted supplementary evaluations using two additional, independent LLM evaluators: Gemini-1.5-pro555https://deepmind.google/technologies/gemini/pro/ and Deepseek-chat666https://www.deepseek.com/. As illustrated in Figure 5, the final Elo scores across the target models show a high degree of consistency across all three evaluators, indicating that the relative rankings of the target models are largely insensitive to the choice of evaluator, suggesting that our findings are not an artifact of any particular evaluator’s biases or limitations.
Additionally, to examine potential length bias Wei et al. (2024), we analyzed dialogue lengths and turn counts across models (Figure 6). The analysis shows no significant correlation, indicating no observable bias.
| Model | Overall | Zero Shot | Few Shot | Win Counts |
| Tongyi-Finance-14B-Chat | 1156.99 | 1007.52 | 1171.27 | 3614 |
| CFGPT2-7B | 1155.75 | 1125.33 | 1157.93 | 3972 |
| Palmyra-Fin-70B-32k | 1128.25 | 1028.18 | 1143.66 | 3634 |
| GPT-4o | 1117.68 | 979.85 | 1120.89 | 3040 |
| Llama-3.1-8B-Instruct | 1046.87 | 1062.18 | 1051.32 | 3294 |
| Internlm2.5-7b-chat | 995.85 | 1009.78 | 1000.52 | 2964 |
| Llama3-XuanYuan3-70B-Chat | 913.48 | 934.51 | 911.59 | 2050 |
| Llama-3.1-70B-Instruct | 912.26 | 986.77 | 906.80 | 2196 |
| GPT-4o-mini | 901.75 | 943.81 | 908.92 | 2326 |
| Qwen2.5-14B-Instruct | 855.82 | 974.27 | 840.05 | 1774 |
| Qwen2.5-7B-Instruct | 814.48 | 946.45 | 786.28 | 1312 |
5.4 Human Preference Alignment
To further validate that our results aligned with actual preferences of users, we also conducted a human preference evaluation involving 15 financial professionals and students, with each assigned 10 pre-existing results to conduct the human preference alignment phase. After manually reassessing the outputs, we updated the Elo scores and compared them with our model’s predictions. To quantify the similarity between the model’s results and human evaluations, we used the Pearson correlation coefficient:
where and represent the Elo scores from participants and model predictions, respectively. The analysis revealed a clear positive correlation, with a calculated Pearson correlation of , as shown in Figure 7, indicating that the model’s performance aligns well with human preferences.
5.5 Case Study
Previous research has demonstrated the significant influence of scaling laws on model performance Kaplan et al. (2020); Ruan et al. (2024). As shown in Figure 8, our results also show a similar trend. LLMs with larger parameters generally outperform smaller ones within the same series, also LLMs from the same backbone model have better results after being trained on financial corpus. However, Llama3.1 appears to be an outlier in this pattern. As shown in Figure 9, we illustrate a result of the Summarization task that highlights this phenomenon. In many of the data points, we observe that the 70B model tends to generate longer, more verbose outputs compared to smaller models. This aligns with the conclusions of Chiang and yi Lee (2024), where larger models are prone to over-reasoning, generating lengthy and unnecessary responses to questions. In our evaluation framework, where clear scoring criteria have been established, these redundant outputs significantly lower the model’s performance.
6 Conclusion
In this paper, we introduced the UCFE Benchmark, a framework designed to evaluate user-AI interactions in the financial domain using the LLM-as-Judge methodology. This framework enables direct comparisons of model performance with human expert preferences while addressing potential biases. Our findings demonstrate that LLMs trained on domain-specific financial texts show notable improvements in understanding complex financial concepts and accurately interpreting user intent. Notably, mid-sized models (7B to 14B parameters) performed particularly well, striking an effective balance between computational efficiency and domain-specific expertise without the excessive overhead of larger models. These results emphasize the importance of optimizing LLMs not only for performance but also for resource efficiency, making them more viable for real-world financial applications. Additionally, the user-centric design of our benchmark highlights the critical role of aligning AI systems with diverse user needs, ensuring that LLMs deliver practical, contextually relevant solutions in finance. This approach lays the foundation for more reliable and scalable AI-driven innovations in the financial industry.
Limitation
The limitations of our work can be summarized as follows:
Coverage of Financial Tasks: The financial domain encompasses a wide range of complex tasks and scenarios, from regulatory compliance to dynamic market analysis. While the UCFE Benchmark includes several representative tasks, the diversity and volume of data points may not be sufficient to fully capture all real-world financial applications. This limitation restricts the benchmark’s ability to comprehensively assess LLM performance across the entire spectrum of financial use cases.
Human Preference Bias: The evaluation framework relies on human preferences to assess model performance, which introduces potential biases. Given the limited number of evaluators and the relatively narrow range of professional backgrounds represented, the results may not fully reflect the diverse needs and preferences of the broader financial community. Individual biases and subjective judgments could influence the evaluation, potentially skewing the assessment of LLM effectiveness in real-world financial tasks.
Use of Historical Data: The benchmark relies primarily on historical financial data for task evaluation. While this data is useful for assessing LLM performance in past scenarios, it may not fully capture the evolving and real-time nature of financial markets. This reliance on historical data limits the ability to evaluate how well LLMs can adapt to unforeseen events or respond to rapidly changing market conditions.
Ethical Statements
We do not see our work to have possible harmful outcomes. We follow the ACL ethical guidelines when conducting the research in this paper.
Acknowledgement
This work was supported by the Major Program of the National Fund of Philosophy and Social Science of China (No. 19ZDA105), the Shenzhen Science and Technology Program (JCYJ20220818103001002), the Shenzhen Doctoral Startup Funding (RCBS20221008093330065), the Tianyuan Fund for Mathematics of the National Natural Science Foundation of China (NSFC) (12326608), the Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (Grant No. ZDSYS20230626091302006), and the Shenzhen Stability Science Program 2023.
Appendix
Appendix A Questionnaire
The questionnaire of our survey is shown in Figure A-1.
Appendix B UCFE Dataset Information
B.1 Geographical Distribution
Figure B-1 shows the geographical distribution of our previous survey. Among our 804 participants, 62.9% of them are from China, 35.9% from the USA, and 1.2% from other regions. This highlights the dominance of responses from China and the USA.
B.2 Results of the survey
Figure B-2 and Figure B-3 report the primary financial information source and the results of whether users prefer generation answers or predefined options, which demonstrates the diversity of our benchmark contributors.
B.3 Detailed Information of Each Task
B.3.1 Zero-shot Tasks
Similar to existing benchmarks, the zero-shot tasks in the UCFE benchmark require models to handle new financial problems without any prior examples. These tasks assess the models’ ability to generalize across different types of financial challenges. The UCFE benchmark includes four zero-shot tasks:
Stock Price Prediction: Predicting future stock prices using historical A-stock statistics is a common task in financial forecasting.
Bearish Information Detection: Identifying whether the information affects the market negatively from sources such as EastMoney, similar to risk detection tasks in other benchmarks.
Financial Indicator Calculation: Computing important financial metrics using standard CPA and CFA formulas, much like quantitative tasks in existing financial benchmarks.
Financial Information Summarization: Summarizing news headlines to capture key insights, a task also present in general NLP benchmarks but adapted to the financial context.
These tasks are designed to reflect real-world financial decision-making scenarios, targeting two broad user groups: the general public and financial professionals. In essence, they encompass a wide range of users, making the benchmark applicable to all types of financial stakeholders.
B.3.2 Few-shot Tasks
The few-shot tasks in the UCFE benchmark involve multi-turn financial interactions, focusing on how models adapt to evolving user input over several rounds. Unlike single-turn tasks in existing benchmarks, these tasks emphasize real-world financial decision-making. We categorize the 13 tasks into the following four main groups:
Analytical and Evaluation Tasks
These tasks require the model to simulate the role of financial analysts, providing detailed insights based on iterative queries. The model must refine its responses as users ask follow-up questions:
Analyst Simulation: Comprehensive analysis of company performances from financial reports and analyst reviews and generate recommendations.
Asset Valuation Reporting: Provide asset valuations using data from EastMoney.
Company Evaluation Reporting: Evaluate company performance using financial reports.
Corporate Operation Analysis: Analyze company operations based on analyst reports.
Risk and Compliance Tasks
These tasks focus on identifying financial risks and ensuring compliance with regulations, where users interact with the model to iteratively refine their analysis:
Credit Risk Evaluation: Assess credit risks based on GPT-4-generated data.
Insider Trading Detection: Identify potential insider trading cases using court records of historical insider trading case reports.
Prospectus Risk Summarization: Summarize risks in prospectuses and inquiry letters, refining insights based on user feedback.
Financial Regulation Consulting: Provide guidance on regulatory compliance and potential punishments using Securities Law.
Strategy and Optimization Tasks
In these tasks, users interact with the model to evaluate and optimize investment strategies. The multi-turn nature allows users to explore different strategies or fine-tune their approach:
Investment Strategy Evaluation: Evaluate effectiveness and summarize investment strategies using data from Seeking Alpha.
Investment Strategy Optimization: Optimize strategies with feedback from multiple rounds of user queries.
Consulting and Summarization Tasks
These tasks involve providing consulting services or summarizing financial information, where users may request additional clarification or focused insights over several interactions:
Financial Knowledge Consulting: Offer explanation on financial terminologies and basic financial knowledge based on sources like Investopedia.
Industry Report Summarization: Summarize industry reports from EastMoney, allowing users to quickly identify key trading insights.
Newshare Evaluation Reporting: Evaluate target price range of newly issued shares, analyze risk and opportunity of the new share based on company overview using data from platforms like stock.us.
Appendix C Human Expert Evaluation
Figure C-1 illustrates the evaluation pipeline. Human experts first assess the existing results, and if the outcomes directly contradict their expertise, they are asked to make another selection. Additionally, we designed a UI interface, shown in Figure C-2, to ensure that participants are unaware of the model names. After all the evaluations, we recalculated the Elo scores based on the samples evaluated by the human experts.
Appendix D More Experiment Results
Figure B-4 highlights each model’s comparative performance, showcasing strengths and weaknesses. And Table D-1 presents the Elo scores across all tasks. The results show that LLMs trained on financial text data demonstrate outstanding performance across various tasks. However, due to limitations in the zero-shot task texts, the models’ performance in those tasks may be affected.
Figure C-3 shows an example dialogue between User Simulator and Tongyi-Finance in the “Financial Knowledge Consulting” Task.
| Model | Tongyi-Finance-14B-Chat | CFGPT2-7B | Palmyra-Fin-70B-32k | GPT-4o | Llama-3.1-8B-Instruct | Internlm2.5-7b-chat | Llama3-XuanYuan3-70B-Chat | Llama-3.1-70B-Instruct | GPT-4o-mini | Qwen2.5-14B-Instruct | Qwen2.5-7B-Instruct |
| Analyst Simulation | 1114.25 | 1188.80 | 1036.15 | 1066.31 | 1027.62 | 1018.49 | 926.93 | 915.40 | 933.63 | 875.52 | 895.78 |
| Asset Valuation Reporting | 1077.59 | 1202.57 | 1142.88 | 1033.36 | 959.37 | 1060.53 | 871.43 | 939.43 | 956.57 | 931.89 | 823.02 |
| Company Evaluation Reporting | 1068.18 | 1194.01 | 1000.03 | 1038.43 | 1027.47 | 1047.91 | 886.98 | 917.06 | 1043.14 | 891.65 | 883.52 |
| Corporate Operation Analysis | 1178.15 | 1053.48 | 1097.44 | 1076.33 | 1020.20 | 909.57 | 980.96 | 1006.09 | 977.73 | 909.79 | 788.69 |
| Credit Risk Evaluation | 1151.50 | 1209.96 | 1121.60 | 1082.85 | 1028.88 | 1039.29 | 992.64 | 956.07 | 888.36 | 806.93 | 719.89 |
| Financial Knowledge Consulting | 1181.90 | 1275.16 | 1130.83 | 1073.20 | 1062.81 | 958.14 | 920.05 | 902.51 | 875.90 | 869.82 | 747.07 |
| Financial Regulation Consulting | 1070.82 | 1159.60 | 1128.95 | 1035.19 | 1034.53 | 1025.16 | 1008.68 | 982.39 | 891.72 | 865.44 | 795.66 |
| Industry Report Summarization | 996.86 | 1035.47 | 1154.38 | 1149.73 | 1050.53 | 1003.94 | 888.71 | 928.58 | 978.90 | 947.57 | 863.84 |
| Insider Trading Detection | 1031.29 | 1036.04 | 1139.13 | 1058.42 | 1180.18 | 1054.34 | 936.75 | 968.44 | 975.41 | 855.87 | 763.09 |
| Investment Strategy Evaluation | 1223.95 | 1173.37 | 1074.47 | 999.73 | 1086.70 | 969.71 | 929.73 | 981.23 | 928.18 | 843.62 | 787.67 |
| Investment Strategy Optimization | 1177.51 | 1140.73 | 1131.13 | 997.38 | 1087.51 | 1035.26 | 961.40 | 981.92 | 898.98 | 829.85 | 756.83 |
| Newshare Evaluation Reporting | 989.39 | 1037.57 | 1056.65 | 1114.74 | 1030.47 | 982.26 | 925.07 | 879.8 | 1052.01 | 1050.42 | 880.39 |
| Prospectus Risk Summarization | 1171.76 | 1123.44 | 1122.57 | 1110.57 | 979.56 | 953.88 | 899.72 | 941.71 | 922.92 | 835.63 | 936.02 |
| Stock Price Prediction | 1018.66 | 981.21 | 983.73 | 1012.44 | 967.88 | 1003.67 | 985.76 | 1000.01 | 1024.75 | 1011.70 | 1010.08 |
| Negative Information Detection | 998.72 | 1113.40 | 998.85 | 978.35 | 1007.33 | 978.85 | 970.53 | 980.32 | 980.59 | 1009.53 | 983.24 |
| Financial Indicator Calculation | 1119.17 | 1180.77 | 1067.67 | 858.10 | 1131.91 | 1061.85 | 972.09 | 928.13 | 861.15 | 897.84 | 919.85 |
| Financial Text Summarization | 974.17 | 1103.51 | 1030.04 | 1022.80 | 1095.63 | 982.84 | 909.35 | 1027.29 | 954.68 | 971.29 | 927.49 |
| Overall | 1156.99 | 1155.75 | 1128.25 | 1117.68 | 1046.87 | 995.85 | 913.48 | 912.26 | 901.75 | 855.82 | 814.48 |
Appendix E Prompt
E.1 Zero Shot Task
Figures [E-1, E-2, E-3] have shown all the prompts we used for testing and evaluation for Financial Text Summarization.
E.2 Few Shot Task
Figures [E-4, E-5, E-6] have shown all the prompts we used for testing and evaluation for Asset Valuation Reporting.